# Sensitivity Analysis on Loss Landscape

### 02 Mar 2024 v1

@@ Abstract
Gradients can be employed for sensitivity analysis. Here, we leverage the advantages of the Loss Landscape to comprehend which independent variables impact the dependent variable. We seek to grasp the loss landscape by utilizing first, second, and third derivatives through automatic differentiation. we know that Spearman’s rank correlation coefficient can detect the monotonic relationship between two variables. However, I have found that second-order gradients, with certain configurations and parameters, provide information that can be visualized similarly to Spearman’s results.In our approach, we incorporate a loss function with an activation function, resulting in a non-linear pattern. Each exploration of the loss landscape through retraining yields new valuable information. Furthermore, the first and third derivatives are also beneficial, as they indicate the extent to which independent variables influence the dependent variable.
@@ 3.3 Navigating through automatic differentiation
Automatic differentiation (AD) is a set of techniques used to compute the derivative of a function specified by a computer program. It exploits the fact that every computer calculation, no matter how complicated, executes a sequence of elementary arithmetic operations and functions. By repeatedly applying the chain rule 2 to these operations, AD can compute partial derivatives of arbitrary order accurately.
@@ 4 Loss Landscape Perspective
Loss landscape is changes as we change the loss function so we stick at more reliable function , so the choice is Mean Squared Error, and loss surface also changes by activation function so as I mentioned above activation function is chosen by experiment , even though as we can imagine them intuitively the surface affected by from multiple parameters such as weights so we have sticking with Xavier glorot initialization 3, the choice of Normal distribution is very important since we want to explore the loss landscape a lot by every time we retrain it , but the uniform distribution will limit the exploration a bit , so we are gonna stick with normal. The structure of neural loss functions and their landscapes significantly influence the generalization of neural networks. also that certain network architectures and training parameters can shape these landscapes and the minimizers found within them.[7]
@@ 5.1 Loss Landscape of Second Order Partial Derivative
In the context of this model setup, where we only have a single column of second-order gradients due to the limitations of calling it a Hessian matrix, now take a function $f(x)$ and its second-order partial derivative 4, for a given function. This derivative provides insights into the curvature of the function with respect to the variable $x$. Specifically, a positive second derivative indicates a local minimum. if we are at a local minimum, the second-order gradient is expected to be positive. This is a crucial observation in understanding the behavior of the neural network during training.
@@ 5.4 Possibility of Gradients in images
I conducted a small experiment using the MNIST dataset. I trained a model using a subset of the dataset and stored the second order gradients. I then analyzed the second order gradient to identify the elements that increased and decreased. The decreasing elements were deemed not useful for prediction. From this, I created a single mask using the gradients and applied it to the MNIST image 5. However, there are several limitations to this approach. For instance, the computational requirements are high and the MNIST dataset is not representative of all image datasets due to its fixed portion. Nonetheless, this was an experimental thought and a step towards understanding the potential of gradients in image processing.

### 05 Mar 2024 v2

@@ Abstract
Gradients can be employed for sensitivity analysis. Here, we leverage the advantages of the Loss Landscape to comprehend which independent variables impact the dependent variable. We seek to grasp the loss landscape by utilizing first, second, and third derivatives through automatic differentiation. we know that Spearman’s rank correlation coefficient can detect the monotonic relationship between two variables. However, I have found that second-order gradients, with certain configurations and parameters, provide information that can be visualized similarly to Spearman results,In this approach, we incorporate a loss function with an activation function, resulting in a non-linear pattern. Each exploration of the loss landscape through retraining yields new valuable information. Furthermore, the first and third derivatives are also beneficial, as they indicate the extent to which independent variables influence the dependent variable.
@@ 3.3 Navigating through automatic differentiation
Automatic differentiation (AD) is a set of techniques used to compute the derivative of a function specified by a computer program. It exploits the fact that every computer calculation, no matter how complicated, executes a sequence of elementary arithmetic operations and functions. By repeatedly applying the chain rule (2) to these operations, AD can compute partial derivatives of arbitrary order accurately.
@@ 4 Loss Landscape Perspective
Loss landscape is changes as we change the loss function so we stick at more reliable function , so the choice is Mean Squared Error, and loss surface also changes by activation function so as I mentioned above activation function is chosen by experiment , even though as we can imagine them intuitively the surface affected by from multiple parameters such as weights so we have sticking with Xavier glorot initialization (3), the choice of Normal distribution is very important since we want to explore the loss landscape a lot by every time we retrain it , but the uniform distribution will limit the exploration a bit , so we are gonna stick with normal. The structure of neural loss functions and their landscapes significantly influence the generalization of neural networks. also that certain network architectures and training parameters can shape these landscapes and the minimizers found within them.[7]
@@ 5.1 Loss Landscape of Second Order Partial Derivative
In the context of this model setup, where we only have a single column of second-order gradients due to the limitations of calling it a Hessian matrix, now take a function $f(x)$ and its second-order partial derivative (4), for a given function. This derivative provides insights into the curvature of the function with respect to the variable $x$. Specifically, a positive second derivative indicates a local minimum. if we are at a local minimum, the second-order gradient is expected to be positive. This is a crucial observation in understanding the behavior of the neural network during training.
@@ 5.4 Possibility of Gradients in images
I conducted a small experiment using the MNIST dataset. I trained a model using a subset of the dataset and stored the second order gradients. I then analyzed the second order gradient to identify the elements that increased and decreased. The decreasing elements were deemed not useful for prediction. From this, I created a single mask using the gradients and applied it to the MNIST image Figure 5. However, there are several limitations to this approach. For instance, the computational requirements are high and the MNIST dataset is not representative of all image datasets due to its fixed portion. Nonetheless, this was an experimental thought and a step towards understanding the potential of gradients in image processing.